# A Quadratic Lower Bound for Stable Roommates Solvability

### v1

@@ Abstract
In their seminal work on the Stable Marriage Problem (SM), Gale and Shapley introduced a generalization of SM referred to as the Stable Roommates Problem (SR). An instance of SR consists of a set of $2n$ agents, and each agent has preferences in the form of a ranked list of all other agents. The goal is to find a one-to-one matching between the agents that is stable in the sense that no pair of agents have a mutual incentive to deviate from the matching. Unlike the (bipartite) stable marriage problem, in SR, stable matchings need not exist. Irving devised an algorithm that finds a stable matching or reports that none exists in $O(n^{2})$ time. In their influential 1989 text, Gusfield and Irving posed the question of whether $\Omega(n^{2})$ time is required for SR solvability—the task of determining if an SR instance admits a stable matching.
In this paper we provide an affirmative answer to Gusfield and Irving’s question. We show that any (randomized) algorithm that determines SR solvability requires $\Omega(n^{2})$ adaptive Boolean queries to the agents’ preferences (in expectation). Our argument follows from a reduction from the communication complexity of the set disjointness function. The query lower bound implies quadratic time lower bounds for Turing machines, and memory access lower bounds for random access machines. Thus, we establish that Irving’s algorithm is optimal (up to a logarithmic factor) in a very strong sense.
@@ 1 Introduction
Since its introduction by Gale and Shapley, SR has been the subject of a large volume of theoretical work and has found many applications. As the problem’s name suggests, SR naturally models situations in which “peers” should be matched with one another, such as identifying roommates in housing searches [17]. Algorithms for SR have been employed to facilitate kidney exchange markets for organ donation in several countries [1, 2, 4]. SR algorithms have also been employed for peer-to-peer file sharing networks [14, 16], and forming pairs of players in chess tournaments [12]. We refer the readers to [15] for a thorough overview of SR and applications.
Unlike the bipartite SM—where stable marriages are guarantee to exist for all preferences—Gale and Shapley observed that SR instances need not admit stable matchings (see Section 2.1.1 for an explicit example). The question of whether or not a given instance admits a stable matching is known as the SR solvability problem. Gale and Shapley [6] and subsequently Knuth [11] asked whether there exists an efficient algorithm for SR solvability (and finding a stable matching if one exists), where Knuth suggested the problem might be NP-complete. In 1985, Irving [9] devised an efficient algorithm for SR solvability. Irving’s algorithm finds a stable matching or reports that none exists in $O(n^{2})$ time on instances with $2n$ agents. [^1] Observe that since the input consists of $2n$ lists each of length $12n-$ 1, $O(n^{2})$ is linear in the input size. In their influential 1989 text on stable matching problems, Gusfield and Irving [8] listed finding a lower bound (or $o(n^{2})$ time algorithm) for SR solvability as one of 12 open questions for future research. This question was again listed as an open problem in Manlove’s comprehensive 2012 text on algorithmic aspects of stable matchings [15]. To our knowledge, the question has remained open since.
@@ 1.1 Our Contributions
Our main result is to prove an $\Omega(n^{2})$ lower bound for any (randomized or deterministic) algorithm solving SR solvability. While Gusfield and Irving’s question is phrased in terms of running time lower bounds for a particular representation of the agents’ preferences, we give a more general result that bounds the number of Boolean queries necessary to determine SR solvability.
@@
[Informal, c.f. Theorem 4] Any algorithm that determines SR solvability requires $\Omega(n^{2})$ Boolean queries to the agents’ preferences for instances with $2n$ agents. This lower bound applies to randomized protocols (in expectation) and allows for arbitrary Boolean queries made to individual agents’ preferences as well as queries made to predetermined batches of agents (i.e., queries that involve more than one agent’s preferences).
One advantage of phrasing our lower bound in terms of (arbitrary) Boolean queries is that the lower bound extends to many models of computation simultaneously and is agnostic to the representation of the preferences.
@@
The following lower bounds hold for SR solvability of instances with $2n$ agents:
@@
    Any (multi-tape, probabilistic) Turing machine that solves SR solvability requires $\Omega(n^{2})$ time (in expectation).
@@
    Any (randomized) random access machine (RAM) with word size $O(\log n)$ bits that solves SR solvability requires $\Omega(n^{2}/\log n)$ memory accesses (in expectation).
@@
Another interpretation of our lower bound is for mechanisms that determine SR solvability by eliciting agents’ implicit preferences. In practice, agents’ preferences may not be known explicitly, even to the agents themselves. [^2] For example, consider the eponymous application of determining suitable roommates for housing. The question of determining if $a$ prefers potential roommate $b$ to $c$ may rely on knowing a significant amount of information about $b$ and $c$ that is not initially known to $a$. In cases of assigning university housing (where potential roommates may not know each other beforehand) responses to questionnaires are often used as a proxy for explicit preferences, with the hope that responses to the questions elicit sufficient information about the “true” preferences to determine roommate compatibility. Thus, it is desirable to design a mechanism (such as questionnaires) that elicits sufficient information about the agents’ preferences to determine SR solvability. Theorem 1 implies that any such mechanism requires $\Omega(n^{2})$ Boolean queries. Thus, determining SR solvability is essentially as hard as learning the agents’ preferences.
In order to prove Theorem 1, we employ a reduction from the two-party communication complexity of the set disjointness function, $\operatorname{disj}$. This technique was previously applied by Gonczarowski et al. [7] to obtain lower bounds for the (bipartite) SM. Given the framework of query lower bounds via communication complexity [3, 5], our construction and argument are quite simple. The basic idea is to define a family of SR instances that correspond to inputs to the disjointness function. That is, given an input $(x,y)$ to $\operatorname{disj}$, we define an SR instance $R(x,y)$ such that $R(x,y)$ admits a stable matching if and only if $\operatorname{disj}(x,y)=1$. We then argue that the correspondence $(x,y)\mapsto R(x,y)$ has the property that any algorithm that determines SR solvability using $q$ queries implies the existence of a communication protocol for $\operatorname{disj}$ using $q$ bits of communication. Our main result follows from well-known communication complexity lower bounds for $\operatorname{disj}$ [10, 18]. We describe this technique more thoroughly in Section 2.2 before defining the reduction formally in Section 3.
@@
We note that our $\Omega(n^{2})$ query lower bound is only tight to Irving’s algorithm [9] up to a logarithmic factor. This is because an SR instance of size $2n$ contains $2n$ lists of $12n-$ 1 agents. If each agent is encoded with $O(\log n)$ bits, the total size of preference lists is $\Theta(n^{2}\log n)$ bits. Irving’s algorithm allows for access to a single entry on a preference list as a unit cost operation, whereas this operation would correspond to $\Theta(\log n)$ Boolean queries in our model. Thus, in our model, Irving’s algorithm uses $\Theta(n^{2}\log n)$ Boolean queries. We leave it as an open question whether the lower bound can be improved to $\Omega(n^{2}\log n)$ Boolean queries (i.e., a lower bound that is truly linear in the instance size) or if the logarithmic factor in the running time can be improved.
@@ 1.2 Related Work
As already mentioned, SR was introduced by Gale and Shapley [6] who showed that SR instances need not admit stable matchings. Knuth [11] explicitly posed the question of whether there is an efficient algorithm to find a stable matching or report that none exists. Irving [9] devised an algorithm that finds a stable matching or correctly determines that none exists in $O(n^{2})$ time (assuming each preference list entry can be accessed in $O(1)$ time). Gusfield and Irving’s influential book [8] gives a comprehensive overview of early algorithmic work on SR, including structural aspects of the set of stable matchings for an SR instance. Additionally, [8] listed two questions related to efficiently determining SR solvability. The first question asked if it is possible to generate a “succinct certificate” for non-solvability of an SR instance. This question was positively answered by Tan [19] who demonstrated that a “stable partition” (of size $O(n)$) can be computed in $O(n^{2})$ time. Given the partition, solvability of an SR instance can be checked in $O(n)$ time. Gusfield and Irving’s second question asked if it is possible to determine SR solvability in $o(n^{2})$ time directly. Our lower bound gives a negative answer to this question. Given Tan’s result, our lower bound also implies that finding a stable partition requires $\Omega(n^{2})$ time. We refer the reader to the text of Manlove [15] for an account of more recent developments related to SR.
@@ 2.1 The Stable Roommates Problem
An SR instance is solvable if it admits a stable matching. The SR Solvability problem is the problem of determining if a given SR instance is solvable.
An efficient ($O(n^{2})$ time) algorithm for solving SR Solvability was proposed by Irving in [9]. We will not require a description of the full algorithm, but our analysis will rely on an initial processing of the preference lists from Irving’s algorithm. Our terminology and exposition of this follow that in Gusfield and Irving’s text [8, Chapter 4].
@@
Lemma 2.1 (Irving [9], c.f. [8, Section 4.2.2]).
Consider the preference table resulting from an execution of Algorithm 1 on an SR instance. Then
@@
    The resulting preference table is independent of the order in which free agents propose ([8, Lemma 4.2.1]).
@@
    Any pair $\left\{a,b\right\}$ that is not contained in the table is not contained in any stable matching. In particular, if any preference list in the table is empty, then the instance is not solvable ([8, Lemma 4.2.3])
@@
    If $M$ is a matching such that all pairs in $M$ appear in the preference table, then no pair $\left\{a,b\right\}$ not appearing the preference table can block $M$. In particular, if the preference table consists only of a matching $M$, $M$ is the unique stable matching for the instance (c.f., [8, Lemmas 4.2.1 and 4.2.4])
@@ 2.1.1 An Unsolvable SR Instance
To see why this instance is not solvable, observe that agents $a$, $b$, and $c$ form a “directed triangle” where they all rank one another in the top two, but the most preferred agents form a cycle $a\to b\to c\to a$. Thus in any matching $M$, $d$’s partner in $M$ will form a blocking pair with one of the other two matched agents. For example, if $\left\{a,d\right\}\in M$, then $\left\{a,b\right\}$ forms a blocking pair.
@@ 2.2 Communication Complexity and Embeddings
Under a generalization of these conditions, Eden and Rosenbaum [5] showed that communication lower bounds imply query lower bounds. Specializing their result to our setting gives the following result.
@@
Suppose there exists an embedding from $\operatorname{disj}$ instances of size $N$ to SR solvability instances of size $n$ satisfying the conditions above. Then any randomized or deterministic protocol for SR solvability for instances of size $n$ must use $\Omega(N)$ queries in expectation.
We elaborate on what the “allowed queries” for our embedding are. In general one cannot allow for arbitrary queries made to arbitrary sets of agents. Indeed, this would allow queries such as “Is the instance solvable?” which trivially solves SR solvability with a single query. At the other extreme, one could restrict queries of particular forms such as “Does agent $a$ prefer agent $b$ to agent $c$?” Our lower bound holds for arbitrary Boolean queries made to any fixed subsets of agents we call batches, so long as each batch contains at most $n/4$ agents. This allows for arbitrary queries to individual agents’ preferences, but also queries whose responses depend on multiple agents within a batch, such as “Does any agent in batch $A$ prefer agent $b$ to agent $c$?”
@@ 3 The Lower Bound
Figure 1: An illustration of the embedding of disjointness for $N=3\times 3$ and $n=3$. This instance corresponds to $x_{1,1}=x_{1,2}=x_{3,2}=1$, while the remaining values of $x_{ij}=0$, and $y_{2,2}=y_{3,3}=1$ with the remaining values of $y_{ij}=0$. Thus, $\operatorname{disj}(x,y)=1.$ The edges between agents are colored by agent preferences: the dark blue edges from the $a_{i}$ correspond to their most preferred partners (preferred above $c_{i}$). Similarly, the dark red edges from the $b_{j}$ correspond to their most preferred edges. The light blue edges from the $c_{i}$ and pink edges from the $d_{j}$ correspond to those agent’s most preferred partners. Other possible partners are not depicted. Sub-figure (a) represents the remaining pairs in the preference table before the first rounds of proposals, while (b), (c), and (d) depict the remaining pairs after each round of proposals. The final figure depicts the unique stable matching for this instance.
@@
To prove the proposition, we consider an execution of the Phase 1 algorithm, Algorithm 1. We first consider the case where $\operatorname{disj}(x,y)=1$ (case 1 of the proposition). Consider an execution of Algorithm 1 with the agents proposing in the following order. We illustrate an example of the execution of such an instance with $n=3$ in Figure 1.
@@
    Agents in $D$ propose. Specifically, each $d_{j}\in D$ proposes to $b_{j}$. When $b_{j}$ receives the proposal from $d_{j}$, only $c_{i}$s with $y_{ij}=1$ and $a_{i}$s with $y_{ij}=1$ remain on $d_{j}$’s preference list. Since $\operatorname{disj}(x,y)=1$, if $y_{ij}=1$, then $x_{ij}=0$. Therefore, all remaining pairs of the form $\left\{a_{i},b_{j}\right\}$ are removed from the preference table.
@@
We now state and prove our main lower bound for the stable roommates problem. In the formal statement, we allow for any procedure that performs arbitrary Boolean queries to the agents’ preference lists, and queries can even be made to “batches” of agents—i.e., a fixed partition of the agents—so long as no batch contains more than $n/2$ agents. This allows queries of the form, for example, “Does any agent in set $A$ prefer agent $b$ to $b^{\prime}$?” Specifically, our argument holds for any Boolean query that does not involve agents from both sets $A$ and $B$ in the construction above. An upper bound on the batch size is necessary allowing a batch size of $n$ would allow for the query “Does $S$ admit a stable matching?” as a single query.
@@
Any randomized (or deterministic) mechanism that determines SR solvability on instances with $n$ agents using adaptive Boolean query access to the agents’ preferences requires $\Omega(n^{2})$ Boolean queries in expectation to the agents’ preferences. This bound applies even if queries can be made to fixed batches of agents of size up to $n/2$.
@@
Let $\mathcal{A}$ be any algorithm that determines SR solvability using $q$ queries for SR instances of size $n$. Then we can use $\mathcal{A}$ to define a two party communication protocol for $\operatorname{disj}$ using the embedding of Proposition 3.1 as follows. Suppose Alice and Bob hold inputs $x,y\in\left\{0,1\right\}^{n\times n}$, respectively, which are promised to be either disjoint or uniquely intersecting. Note that the set disjointness instance has size $N=n^{2}$. Alice forms preferences of agents in the set $A$ as in the embedding, and Bob does the same for $B$. Thus, the instance has $4n$ agents in total. Note that the preferences of agents in $C$ and $D$ are independent of $x$ and $y$, hence they are known to both Alice and Bob. Alice and Bob then simulate the algorithm $\mathcal{A}$: the response to any query that $\mathcal{A}$ makes to agents in $A$ can be computed by Alice, who then sends the Boolean result to Bob as a single bit. Similarly, the response to any query made to $B$ can be computed by Bob, who then sends the response to Alice. When $\mathcal{A}$ terminates after $q$ queries, both Alice and Bob can compute the output of $\mathcal{A}$ from the communication transcript, hence they both know whether or not the $R(x,y)$ is solvable. By Proposition 3.1, this output determines $\operatorname{disj}(x,y)$.
@@
We observe that the arguments above make no reference to the representation of the input to the problem, and the query lower bound argument holds for any Boolean queries made to the input so long as a single query’s response does not depend on both agents from both $A$ and $B$ preference lists. Following an arbitrary preprocessing of the agents’ preferences where agents in $A$ are processed independently from agents in $B$, each bit of the resulting encoding is determined by either preferences in $A\cup C\cup D$ or preferences in $B\cup C\cup D$. In the former case, reading a single bit of the processed input can be simulated with a single Boolean query to $A$ (as preferences in $C$ and $D$ are fixed) in the former case and $B$ in the latter case. Thus, the lower bounds apply to preprocessed inputs as well. ∎

### v2

@@ Abstract
In their seminal work on the Stable Marriage Problem (SM), Gale and Shapley introduced a generalization of SM referred to as the Stable Roommates Problem (SR). An instance of SR consists of a set of $2n$ agents, and each agent has preferences in the form of a ranked list of all other agents. The goal is to find a one-to-one matching between the agents that is stable in the sense that no pair of agents have a mutual incentive to deviate from the matching. Unlike the (bipartite) stable marriage problem, in SR, stable matchings need not exist. Irving devised an algorithm that finds a stable matching or reports that none exists in $O(n^{2})$ time. In their influential 1989 text, Gusfield and Irving posed the question of whether $\Omega(n^{2})$ time is required for SR solvability—the task of deciding if an SR instance admits a stable matching.
In this paper we provide an affirmative answer to Gusfield and Irving’s question. We show that any (randomized) algorithm that decides SR solvability requires $\Omega(n^{2})$ adaptive Boolean queries to the agents’ preferences (in expectation). Our argument follows from a reduction from the communication complexity of the set disjointness function. The query lower bound implies quadratic time lower bounds for Turing machines, and memory access lower bounds for random access machines. Thus, we establish that Irving’s algorithm is optimal (up to a logarithmic factor) in a very strong sense.
@@ 1 Introduction
Since its introduction by Gale and Shapley, SR has been the subject of a large volume of theoretical work and SR has found many applications. As the problem’s name suggests, SR naturally models situations in which “peers” should be matched with one another, such as identifying roommates in housing searches [19]. Variants of SR can be used to model pair-wise kidney exchanges for organ donation [21]. The general study of this problem has led to the development of algorithms to facilitate kidney exchange markets for organ donation in several countries [1, 2, 5]. SR algorithms have also been employed for peer-to-peer file sharing networks [15, 17], and forming pairs of players in chess tournaments [13]. We refer the readers to [16] for a thorough overview of SR and applications.
Unlike the bipartite SM—where stable marriages are guaranteed to exist for all preferences—Gale and Shapley observed that SR instances need not admit stable matchings (see Section 2.1.1 for an explicit example). The question of whether or not a given instance admits a stable matching is known as the SR solvability problem. Gale and Shapley [7] and subsequently Knuth [12] asked whether there exists an efficient algorithm for SR solvability (and finding a stable matching if one exists), where Knuth suggested the problem might be NP-complete. In 1985, Irving [10] devised an efficient algorithm for SR solvability. Irving’s algorithm finds a stable matching or reports that none exists in $O(n^{2})$ time on instances with $2n$ agents. [^1] Observe that since the input consists of $2n$ lists each of length $12n-$ 1, $O(n^{2})$ is linear in the input size. In 1990, Ng and Hirschberg [18] showed that $\Omega(n^{2})$ time is necessary to find a stable marriage (in the bipartite stable marriage problem). Their result implies the same lower bound holds for finding a stable matching for SR, but the lower bound does not apply to the decision problem of determining whether or not an SR instance admits a stable matching. In their influential 1989 text on stable matching problems, Gusfield and Irving [9] listed finding a lower bound (or $o(n^{2})$ time algorithm) for deciding SR solvability as one of 12 open questions for future research. This question was again listed as an open problem in Manlove’s comprehensive 2012 text on algorithmic aspects of stable matchings [16]. To our knowledge, the question has remained open since.
@@ 1.1 Our Contributions
Our main result is to prove an $\Omega(n^{2})$ lower bound for any (randomized or deterministic) algorithm deciding SR solvability. While Gusfield and Irving’s question is phrased in terms of running time lower bounds for a particular representation of the agents’ preferences, we give a more general result that bounds the number of Boolean queries necessary to decide SR solvability.
@@
[Informal, c.f. Theorem 4] Any algorithm that decides SR solvability requires $\Omega(n^{2})$ Boolean queries to the agents’ preferences for instances with $2n$ agents. This lower bound applies to randomized protocols (in expectation) and allows for arbitrary Boolean queries made to individual agents’ preferences as well as queries made to predetermined batches of agents (i.e., queries that involve more than one agent’s preferences).
One advantage of phrasing our lower bound in terms of (arbitrary) Boolean queries is that the lower bound extends to many models of computation simultaneously and it is agnostic to the representation of the preferences.
@@
The following lower bounds hold for deciding SR solvability of instances with $2n$ agents:
@@
    Any (multi-tape, probabilistic) Turing machine that decides SR solvability requires $\Omega(n^{2})$ time (in expectation).
@@
    Any (randomized) random access machine (RAM) with word size $O(\log n)$ bits that decides SR solvability requires $\Omega(n^{2}/\log n)$ memory accesses (in expectation).
@@
Another interpretation of our lower bound is for mechanisms that decide SR solvability by eliciting agents’ implicit preferences. In practice, agents’ preferences may not be known explicitly, even to the agents themselves. [^2] For example, consider the eponymous application of determining suitable roommates for housing. The question of determining if $a$ prefers potential roommate $b$ to $c$ may rely on knowing a significant amount of information about $b$ and $c$ that is not initially known to $a$. In cases of assigning university housing (where potential roommates may not know each other beforehand) responses to questionnaires are often used as a proxy for explicit preferences, with the hope that responses to the questions elicit sufficient information about the “true” preferences to determine roommate compatibility. Thus, it is desirable to design a mechanism (such as questionnaires) that elicits sufficient information about the agents’ preferences to determine SR solvability. Theorem 1 implies that any such mechanism requires $\Omega(n^{2})$ Boolean queries. Thus, determining SR solvability is essentially as hard as learning the agents’ preferences.
In order to prove Theorem 1, we employ a reduction from the two-party communication complexity of the set disjointness function, $\operatorname{disj}$. This technique was previously applied by Gonczarowski et al. [8] to obtain lower bounds for the (bipartite) SM. Given the framework of query lower bounds via communication complexity [3, 6], our construction and argument are quite simple. The basic idea is to define a family of SR instances that correspond to inputs to the disjointness function. That is, given an input $(x,y)$ to $\operatorname{disj}$, we define an SR instance $R(x,y)$ such that $R(x,y)$ admits a stable matching if and only if $\operatorname{disj}(x,y)=1$. We then argue that the correspondence $(x,y)\mapsto R(x,y)$ has the property that any algorithm that decides SR solvability using $q$ queries implies the existence of a communication protocol for $\operatorname{disj}$ using $q$ bits of communication. Our main result follows from well-known communication complexity lower bounds for $\operatorname{disj}$ [11, 20]. We describe this technique more thoroughly in Section 2.2 before defining the reduction formally in Section 3.
@@
Our $\Omega(n^{2})$ query lower bound is only tight to Irving’s algorithm [10] up to a logarithmic factor. This is because an SR instance of size $2n$ contains $2n$ lists of $12n-$ 1 agents. If each agent’s identity is encoded with $\Theta(\log n)$ bits, the total size of preference lists is $\Theta(n^{2}\log n)$ bits. Irving’s algorithm allows for access to a single entry on a preference list as a unit cost operation, whereas this operation would correspond to $\Theta(\log n)$ Boolean queries in our model. Thus, in our model, Irving’s algorithm uses $\Theta(n^{2}\log n)$ Boolean queries. We leave it as an open question whether the lower bound can be improved to $\Omega(n^{2}\log n)$ Boolean queries (i.e., a lower bound that is truly linear in the instance size) or if the logarithmic factor in the running time can be improved.
@@ 1.2 Related Work
SR was introduced by Gale and Shapley [7] who showed that SR instances need not admit stable matchings. Knuth [12] explicitly posed the question of whether there is an efficient algorithm to find a stable matching or report that none exists. Irving [10] devised an algorithm that finds a stable matching or correctly determines that none exists in $O(n^{2})$ time (assuming each preference list entry can be accessed in $O(1)$ time). Gusfield and Irving’s influential book [9] gives a comprehensive overview of early algorithmic work on SR, including structural aspects of the set of stable matchings for an SR instance. Additionally, [9] listed two questions related to efficiently deciding SR solvability. The first question asked if it is possible to generate a “succinct certificate” for non-solvability of an SR instance. This question was positively answered by Tan [23] who demonstrated that a “stable partition” (of size $O(n)$) can be computed in $O(n^{2})$ time. Given the partition, solvability of an SR instance can be verified in $O(n)$ time. Gusfield and Irving’s second question asked if it is possible to decide SR solvability in $o(n^{2})$ time directly. Our lower bound gives a negative answer to this question. Given Tan’s result, our lower bound also implies that finding a stable partition requires $\Omega(n^{2})$ time. Several previous works [18, 4, 22, 8] proved $\Omega(n^{2})$ lower bounds for finding stable marriages in SM (and a fortiori for finding stable matchings in SR), but none of these results imply lower bounds for the decision problem of SR solvability. We refer the reader to the text of Manlove [16] for an account of more recent developments related to SR.
@@ 2.1 The Stable Roommates Problem
An SR instance is solvable if it admits a stable matching. The SR Solvability problem is the problem of deciding if a given SR instance is solvable.
An efficient ($O(n^{2})$ time) algorithm for solving SR Solvability was proposed by Irving in [10]. We will not require a description of the full algorithm, but our analysis will rely on an initial processing of the preference lists from Irving’s algorithm. Our terminology and exposition of Irving’s algorithm follow Gusfield and Irving’s text [9, Chapter 4].
@@
Lemma 2.1 (Irving [10], c.f. [9, Section 4.2.2]).
Consider the preference table resulting from an execution of Algorithm 1 on an SR instance. Then:
@@
    The resulting preference table is independent of the order in which free agents propose ([9, Lemma 4.2.1]).
@@
    Any pair $\left\{a,b\right\}$ that is not contained in the table is not contained in any stable matching. In particular, if any preference list in the table is empty, then the instance is not solvable ([9, Lemma 4.2.3])
@@
    If $M$ is a matching such that all pairs in $M$ appear in the preference table, then no pair $\left\{a,b\right\}$ not appearing the preference table can block $M$. In particular, if the preference table consists only of a matching $M$, $M$ is the unique stable matching for the instance (c.f., [9, Lemmas 4.2.1 and 4.2.4])
@@ 2.1.1 An Unsolvable SR Instance
To see why this instance is not solvable, observe that agents $a$, $b$, and $c$ form a “directed triangle” where they all rank one another in the top two, but the most preferred agents form a cycle $a\to b\to c\to a$. Thus in any matching $M$, $d$’s partner in $M$ will form a blocking pair with one of the other two matched agents. For example, if $\left\{a,d\right\}\in M$, then $\left\{a,c\right\}$ forms a blocking pair.
@@ 2.2 Communication Complexity and Embeddings
Under a generalization of these conditions, Eden and Rosenbaum [6] showed that communication lower bounds imply query lower bounds. We emphasize that this technique yields adaptive query lower bounds: future queries are allowed to depend on the responses to previous queries. Specializing their result to our setting gives the following result.
@@
Suppose there exists an embedding from $\operatorname{disj}$ instances of size $N$ to SR solvability instances of size $n$ satisfying the conditions above. Then any randomized or deterministic protocol for SR solvability for instances of size $n$ must use $\Omega(N)$ adaptive queries in expectation.
We elaborate on what the “allowed queries” for our embedding are. In general one cannot allow for arbitrary queries made to arbitrary sets of agents. Indeed, this would allow queries such as “Is the instance solvable?” which trivially solves SR solvability with a single query. At the other extreme, one could restrict queries of particular forms such as “Does agent $a$ prefer agent $b$ to agent $c$?” Our lower bound holds for arbitrary Boolean queries made to any fixed subsets of agents we call batches, so long as each batch contains at most $n/4$ agents. This allows for arbitrary Boolean queries to individual agents’ preferences, but also queries whose responses depend on multiple agents within a batch, such as “Does any agent in batch $A$ prefer agent $b$ to agent $c$?”
@@ 3 The Lower Bound
Figure 1: An illustration of the embedding of disjointness for $N=3\times 3$ and $n=3$. This instance corresponds to $x_{1,1}=x_{1,2}=x_{3,2}=1$, while the remaining values of $x_{ij}=0$, and $y_{2,2}=y_{3,3}=1$ with the remaining values of $y_{ij}=0$. Thus, $\operatorname{disj}(x,y)=1.$ The edges between agents are colored by agent preferences: the dark blue edges from the $a_{i}$ correspond to their most preferred partners (preferred above $c_{i}$). Similarly, the dark red edges from the $b_{j}$ correspond to their most preferred edges. The light blue edges from the $c_{i}$ and violet edges from the $d_{j}$ correspond to those agent’s most preferred partners. Other possible partners are not depicted. Sub-figure (a) represents the remaining pairs in the preference table before the first rounds of proposals, while (b), (c), and (d) depict the remaining pairs after each round of proposals. The final figure depicts the unique stable matching for this instance.
@@
To prove the proposition, we consider an execution of the Phase 1 algorithm, Algorithm 1. We first consider the case where $\operatorname{disj}(x,y)=1$ (case 1 of the proposition). Consider an execution of Algorithm 1 with the agents proposing in the following order. By Lemma 2.1, the resulting preference table does not depend on the order of the proposals. We illustrate an example of the execution of such an instance with $n=3$ in Figure 1.
@@
    Agents in $D$ propose. Specifically, each $d_{j}\in D$ proposes to $b_{j}$. When $b_{j}$ receives the proposal from $d_{j}$, only $c_{i}$s with $y_{ij}=1$ remain on $b_{j}$’s preference list, as all $a_{i}$’s with $y_{ij}=1$ were removed after the proposals in Step 1. Since $\operatorname{disj}(x,y)=1$, if $y_{ij}=1$, then $x_{ij}=0$. Therefore, all remaining pairs of the form $\left\{a_{i},b_{j}\right\}$ are removed from the preference table.
@@
We now state and prove our main lower bound for the stable roommates problem. In the formal statement, we allow for any procedure that performs arbitrary adaptive Boolean queries to the agents’ preference lists, and queries can even be made to “batches” of agents—i.e., a fixed partition of the agents—so long as no batch contains more than $n/2$ agents. For example, this allows queries of the form, “Does any agent in set $A$ prefer agent $b$ to $b^{\prime}$?” Specifically, our argument holds for any Boolean query that does not involve agents from both sets $A$ and $B$ in the construction above. An upper bound on the batch size is necessary allowing a batch size of $n$ would allow for the query “Does $S$ admit a stable matching?” as a single query.
@@
Any randomized (or deterministic) mechanism that decides SR solvability on instances with $n$ agents using adaptive Boolean query access to the agents’ preferences requires $\Omega(n^{2})$ Boolean queries in expectation. This bound applies even if queries can be made to fixed batches of agents of size up to $n/2$.
@@
Let $\mathcal{A}$ be any algorithm that determines SR solvability using $q$ queries for SR instances of size $n$. Then we can use $\mathcal{A}$ to define a two party communication protocol for $\operatorname{disj}$ using the embedding of Proposition 3.1 as follows. Suppose Alice and Bob hold inputs $x,y\in\left\{0,1\right\}^{n\times n}$, respectively, which are promised to be either disjoint or uniquely intersecting. Note that the set disjointness instance has size $N=n^{2}$. Alice forms preferences of agents in the set $A$ as in the embedding, and Bob does the same for $B$. Thus, the instance has $4n$ agents in total. Note that the preferences of agents in $C$ and $D$ are independent of $x$ and $y$, hence they are known to both Alice and Bob. Alice and Bob then simulate the algorithm $\mathcal{A}$: the response to any query that $\mathcal{A}$ makes to agents in $A\cup C\cup D$ can be computed by Alice, who then sends the Boolean result to Bob as a single bit. Similarly, the response to any query made to $B\cup C\cup D$ can be computed by Bob, who then sends the response to Alice. When $\mathcal{A}$ terminates after $q$ queries, both Alice and Bob can compute the output of $\mathcal{A}$ from the communication transcript, hence they both know whether or not the $R(x,y)$ is solvable. By Proposition 3.1, this output determines $\operatorname{disj}(x,y)$.
@@
We observe that the arguments above make no reference to the representation of the input to the problem, and the query lower bound argument holds for any Boolean queries made to the input so long as a single query’s response does not depend on the preferences of agents from both $A$ and $B$. Following an arbitrary preprocessing of the agents’ preferences where agents in $A$ are processed independently from agents in $B$, each bit of the resulting encoding is determined by either preferences in $A\cup C\cup D$ or preferences in $B\cup C\cup D$. In the former case, reading a single bit of the processed input can be simulated with a single Boolean query to $A$ (as preferences in $C$ and $D$ are fixed) in the former case and $B$ in the latter case. Thus, the lower bounds apply to preprocessed inputs as well. ∎
Acknowledgements
I am thankful to Christine Cheng for reigniting my interest in the stable roommates problem and to David Manlove for his suggestions to improve this manuscript.